OpenStudy (anonymous):

solve 16(^x+2) = 8(^-x)

OpenStudy (anonymous):

Is this \[16^{x +2}\]?

OpenStudy (anonymous):

Yes, yes it is

OpenStudy (anonymous):

If you notice 16 and 8 are factors of 2 you can convert 16 as a power of 2. For example if one of the numbers were 4 you can make it

OpenStudy (anonymous):

\[2^{2}\]

OpenStudy (anonymous):

Oh so 2^4 will get me 16 and then 2^3 will get me 8 correct?

OpenStudy (anonymous):

Or can i even do it by 8 and have 8^2 and 8

OpenStudy (anonymous):

No, that way won't get you there 8 squared is 64

OpenStudy (anonymous):

oh your right :* silly me haha but i dont get how to get to what x= like how to get X by itself

OpenStudy (anonymous):

Once you have them on even footing 2 square something: Your new exponents would be 4(x+2)=3(-x) This is your new equation. Solve for x.

OpenStudy (anonymous):

Oh ok now all i do is distribute the 4 and 3 so 4x+8=-3x then +3x and stuff and get 7x=-8 is the answer which is x= -7/8 :] right!

OpenStudy (anonymous):

-8/7 =x actually

OpenStudy (anonymous):

Right. That's the method. I haven't checked whether you're right or wrong. You can do that.

OpenStudy (anonymous):

Thank you very much